OpenStudy (anonymous):

find the equation of the tangents through the point (8,-1) to the ellipse 2x^2+5y^2=70

OpenStudy (mortonsalt):

What do you know about the tangent line of a function?

OpenStudy (anonymous):

\[\frac{ xx _{i} }{ a^2 }+\frac{ yy _{i} }{ b^2 }=1\]

OpenStudy (mortonsalt):

Basically, the tangent line is the derivative of the function, yeah?

OpenStudy (anonymous):

the slope yea

OpenStudy (mortonsalt):

If you have slope and a point, you can get the equation by \[y-y_1 = m(x-x_1)\]

OpenStudy (anonymous):

no y=mx+b

OpenStudy (mortonsalt):

No, actually, look at it this way instead: Take the derivative of the function. \[4x(dx) +10y(dx)=0\]

OpenStudy (mortonsalt):

sorry 10y(dy)

OpenStudy (mortonsalt):

Plug in your values for x and y at the point (8,-1).

OpenStudy (solomonzelman):

I don't understand what you did, excuse me...

OpenStudy (mortonsalt):

This is the relation between increments of the variables along the tangent line.

OpenStudy (solomonzelman):

Your function is: \(\large\color{black}{ \displaystyle y=\pm\sqrt{-\frac{2}{5}x^2+14 }}\) And since the point (8,-1) is above the x-axis, you choose the +, so just: \(\large\color{black}{ \displaystyle y=\sqrt{-\frac{2}{5}x^2+14 }}\)

OpenStudy (mortonsalt):

I feel like there's an easier way to do it, though.

OpenStudy (solomonzelman):

YOu need the slope of the function at a point x=8, so find f\('\)(8), and this the slope of your line. The point you already know.

OpenStudy (mortonsalt):

Since the increments are measured from (8, -1), then x-8 and y-(-1).

OpenStudy (mortonsalt):

From this, the equation of the line would be \[32(x-8)-10(y+1)=0\]

OpenStudy (solomonzelman):

Basically: \(\large\color{black}{ \displaystyle t(x)=h'(8)\times(x-8)+f(8)}\) (tangent line to the curve at some point is same as Linearization of a curve at some point) And that is: \(\large\color{black}{ \displaystyle t(x)=h'(8)\times(x-8)+(-1)}\)

OpenStudy (solomonzelman):

by t(x) I am denoting the tangent line.

OpenStudy (mortonsalt):

Don't you think my method would work? I feel like it's a lot simpler.

OpenStudy (solomonzelman):

I don't know, but what I did is a typical method.

OpenStudy (solomonzelman):

If your line ends to be the same the you are corret as well.

OpenStudy (anonymous):

so which is correct way?

OpenStudy (mortonsalt):

You can try either way. Do a check to see if the values correspond, then whichever one does, should be the correct way. :) He did it in the typical way, so maybe you should do his.